# STARC-9: A Large-scale Dataset for Multi-Class Tissue Classification for CRC Histopathology

### v1

@@ 4 Experiments and Results
Dataset description: STARC-9 is a comprehensive multi-tissue classification dataset consisting of 630,000 high-resolution non-overlapping 256×256 pixel tiles extracted from 40x magnification (0.25 micrometers/pixel) WSI. It includes nine clinically relevant tissue types: ADI, LYM, MUS, FCT, MUC, NCS, BLD, TUM, and NOR, capturing diverse, fine-grained tissue morphologies. To facilitate rigorous validation of models trained on STARC-9, two independent validation sets were prepared. (i) STANFORD-CRC-HE-VAL-SMALL contains 18,000 tiles (2,000 per tissue type) obtained from 20 WSI separate from the cases used in STARC-9 and was used for preliminary model testing, yielding 79.59% and 85.9% accuracy for models trained on NCT-CRC-HE-100K (NCT) [7] and HMU-GC-HE-30K (HMU) [8], respectively. This highlighted the need for larger, more diverse training sets, as these models struggled with mixed tissue-type tiles, achieving only 60% overall per-WSI tissue mapping accuracy. (ii) The primary validation set, STANFORD-CRC-HE-VAL-LARGE, includes 54,000 tiles (6,000 per class) from 50 WSI independent from STARC-9 and STANFORD-CRC-HE-VAL-SMALL for performance evaluation. Training (STARC‑9) and validation (STANFORD-CRC-HE-VAL-SMALL and STANFORD-CRC-HE-VAL-LARGE) datasets were drawn from different patients, with no overlap between any of the datasets. Additionally, an external CURATED-TCGA-CRC-HE-VAL-20K dataset was prepared with 20,000 tiles extracted from 30 TCGA-CRC WSI to assess the robustness and generalizability of models trained on STARC-9. During benchmarking, models trained on STARC-9 (630,000 tiles, 9 classes), NCT (100,000 tiles, 9 classes), and HMU (30,000 tiles, 8 classes) were validated on seven overlapping tissue types (ADI, LYM, MUS, MUC, NCS, TUM, and NOR), ensuring fairness in performance evaluation.
@@
Model description: To evaluate the utility of STARC-9, we conducted a series of benchmarking experiments using a diverse set of deep learning models, including baseline CNNs, SOTA transformer models, and pathology-specific foundation models. The objective was to assess classification performance, generalizability, and practical utility compared to models trained (fine-tuned) on publicly available datasets like NCT and HMU. Baseline models included ResNet-50, EfficientNet-B7, KimiaNet, and ViT-base, while SOTA models included DeiT-B, Swin Transformer-Base, and ConvNeXT-Base. Pathology-specific foundation models such as CTransPath, HiPT, Prov-Gigapath, Path-DINO, CONCH, UNI, Virchow, and VIM4Path were also tested to assess their generalizability on diverse tissue morphologies. Each model was trained on the STARC-9, NCT, and HMU datasets with Macenko normalization [21]. All models were fine-tuned with a batch size of 32, learning rate of 0.0001, weight decay of 1e-5, Adam optimization, and data augmentation (horizontal/vertical flips, random rotation, and color jittering) for 10 epochs. Maintaining identical training configurations, including batch size and optimizer settings, across datasets was important for a fair and unbiased model comparison. As our primary objective was to isolate the impact of the dataset on model performance, we kept the training configurations consistent to avoid possible confounding introduced by different hyperparameters being applied to the datasets. Model performance was evaluated using precision, recall, macro F1 score, accuracy, and the number of trainable parameters. STARC-9-trained models consistently outperformed models trained on other datasets, exhibiting better generalizability.
@@
Resource description: All experiments were conducted on the following platforms: (i) a local server with 8x NVIDIA L40S 48GB GPUs, and the Stanford (ii) Carina [22] and (iii) Marlowe high-performance computing platforms [23].
@@
ResNet50 [24] 84.08 87.81 98.92 62.59 85.71 98.64 63.17 86.00 98.78 62.59 85.71 98.64 24 M
EfficientNet-B7 [25] 89.99 88.65 99.11 82.47 87.45 98.80 84.55 87.87 98.95 82.47 84.45 98.80 64 M
ViT-base [26] 92.71 91.57 98.49 84.25 90.29 98.09 87.30 90.87 98.28 84.25 90.29 98.09 86 M
@@
DeiT-B [27] 94.28 90.97 98.99 81.63 90.05 98.65 85.35 90.40 98.81 81.63 90.05 98.65 86 M
Swin Trans-base [28] 90.11 93.17 99.09 79.05 91.88 98.80 82.52 92.46 98.94 79.05 91.88 98.79 87 M
KimiaNet [20] 87.25 88.60 99.03 71.53 86.67 98.72 71.53 87.04 98.87 68.69 86.67 98.72 7M
ConvNeXT-base [29] 91.95 92.09 99.01 82.82 91.07 98.36 85.56 91.50 98.68 82.82 91.07 98.36 88 M
@@
CTransPath [30] 90.11 93.17 99.34 79.05 91.88 99.00 82.52 92.46 99.16 79.05 91.88 99.00 87 M
HiPT [31] 90.92 93.21 98.64 74.51 91.99 98.32 77.41 92.54 98.47 74.51 91.99 98.32 86 M
ProvGigPath [32] 89.43 91.47 98.74 74.18 90.60 98.37 78.40 90.92 98.55 74.18 90.60 98.37 305 M
PathDino [33] 92.93 91.19 98.67 77.35 89.64 98.37 81.71 90.22 98.51 77.35 89.64 98.37 22 M
CONCH [34] 91.53 91.41 98.56 75.69 90.02 98.19 78.08 90.52 98.37 75.69 90.02 98.19 87 M
UNI [35] 94.55 93.03 98.67 80.43 91.80 98.25 84.42 92.36 98.45 80.43 91.80 98.26 88 M
Virchow [36] 92.51 92.35 98.63 79.02 91.23 98.28 82.05 91.69 98.45 79.02 91.23 98.28 305 M
@@ 4.1 Multi-Class Tissue Classification
Table 2 reports the precision, recall, F1‑macro, and accuracy metrics for the top‑performing models (with respect to accuracy in Table 1), when validated on STANFORD-CRC-HE-VAL-SMALL, STANFORD-CRC-HE-VAL-LARGE, and CURATED-TCGA-CRC-HE-VAL-20K. These top-performing models were: ViT‑Base trained on NCT, HiPT trained on HMU, and CTransPath trained on STARC‑9. For STANFORD-CRC-HE-VAL-SMALL, the STARC-9-trained CTransPath model achieved 99.75% precision, 99.73% recall, 99.74% F1- macro, and 99.73% accuracy - significantly higher than the other models, which showed lower recall and F1-macro scores. Similarly, on STARC-9-HE-VAL-LARGE, CTransPath maintained its lead, with 99.34% precision, 99.00% recall, 99.16% F1-macro, and 99.00% accuracy. Even on the more challenging STANFORD-TCGA-CRC-HE-20K set, CTransPath consistently achieved near-perfect precision (99.05%), recall (98.88%), and F1-macro (98.96%), demonstrating excellent generalization and robustness across diverse tissue types. To further evaluate the generalizability of the model trained on STARC-9, we curated a small test set of the seven common tissue classes taken from 10 WSI from the IMP‑CRS10K biopsy/polypectomy dataset [42]. In total, 1,093 image tiles were annotated for model performance validation, in which the STARC‑9-trained model achieved a 95.55% F1‑macro and 96.61% accuracy, consistently outperforming the HMU and NCT–trained models. It would also be interesting to evaluate the performance of the models trained on STARC-9 on the NCT and HMU datasets. However, as noted in Section 2 (and in reference [11]), the publicly available validation sets from NCT and HMU contain a substantial fraction of artifact‐laden or mislabeled tiles, as well as "ambiguous" tiles with more than one tissue type represented within the same tile, despite only a single tissue-type label being assigned to the tile. In order to utilize these two datasets as reliable validation datasets, pathologist re‑verification and correction/refinement of the tile‐level labels would be necessary, which is labor‐intensive and infeasible, given that the original WSI used to generate these two datasets were not publicly available for verification of label accuracy.
@@ 4.2 Tumor Tissue Segmentation
Among the most common downstream applications for multi-class tissue classification is tissue segmentation, especially tumor region segmentation, which allows for the automated identification and cropping of tumor-containing regions for subsequent annotation, ROI selection, diagnosis, and prognostication. This approach also facilitates downstream applications such as MSI [6] and other biomarker status prediction, and supports survival modeling [7] for risk stratification and personalized treatment planning. In this context, we conducted experiments to evaluate the effectiveness of models trained on HMU, NCT, and STARC-9 for tumor segmentation, focusing on their ability to accurately identify tumor regions that are important for clinical decision-making. As there were no publicly available CRC WSI repositories with readily usable, high‑quality TUM masks for a larger scale experiment, and existing weakly‑supervised tools did not provide the precision required for generating tissue segmentation masks, we prepared a test set by enlisting pathologists to manually annotate (ground truth) the TUM region in patches of size 2048×2048 pixels using QuPath [13]. We selected 45 patches (3 per slide) from 15 Stanford WSI and 50 patches (2 per slide) from 25 TCGA-CRC WSI, which were fully independent of our training and validation sets, for a more controlled evaluation of model performance [4]. Some patches contained mixed tissue types in order to evaluate the effectiveness of the trained models. For example, as shown in Figure 4, one region contained predominantly tumor, while the other included a mix of tumor and non-tumor tissue (NCS), providing more challenging segmentation.
@@ Appendix D Additional experiments with recent models
We trained and evaluated recent models such as TransNeXt [43], OverLoCK [44], and Beit-base [45] with masked image modeling on the NCT, HMU, and STARC-9 datasets. The comprehensive results are presented in Table 4. While these recent models demonstrate competitive performance, they did not surpass our best-performing combinations: CTransPath on STARC-9, HiPT on HMU, and ViT-base on NCT (as reported in Table 1. Notably, models trained on STARC-9 consistently achieve superior performance across all architectures, validating our dataset’s quality and diversity.
@@
TransNeXt [43] 59.68 56.57 98.60 73.89 64.22 98.61 110M
OverLoCK [44] 63.75 62.09 97.20 79.00 68.98 97.22 24.3M
Beit-base [45] 59.95 76.48 98.17 74.08 86.17 98.20 86.5M
@@
TransNeXt [43] 59.62 55.05 98.58 73.55 62.78 95.68 110M
OverLoCK [44] 63.69 55.56 98.85 79.29 64.09 98.87 24.3M
Beit-base [45] 62.01 78.11 98.61 77.03 88.40 98.68 86.5M
@@
TransNeXt [43] 58.95 51.92 95.57 72.65 61.70 96.07 110M
OverLoCK [44] 64.47 53.93 95.38 80.25 62.98 95.56 24.3M
Beit-base [45] 57.45 72.57 97.93 71.75 82.11 97.87 86.5M
@@ Appendix G Ablation study
Number of samples per cluster ($m$): Determining the number of clusters ($K$) and $m$ is very challenging, and especially with histopathology-based tiles, it is difficult to set one value for the number of clusters and samples. We tested $m=100$ to $800$, finding that a small $(m\approx 100)$ yielded many tiny, redundant clusters, while a large $(m>800)$ resulted in mixing of tissue types. We found that $m\approx 400$ balanced intra-cluster purity and inter-cluster diversity. Table 6 shows the trade-off between the size of the cluster and the purity (inverse of the Shannon entropy) based on the fixed number of significant variations in TUM and NCS tissue morphology. As $m$ increases, purity decreases (entropy rises). We selected $m=400$, as it yielded substantial morphologic variation while maintaining low tissue-type admixture, representing the best balance for our goal. If deciding $m$ is complex for a dataset, it is recommended that the number of clusters and samples per cluster be set using $K=m=\sqrt{T}$ [46], where $T$ is the number of tiles from a WSI.

### v2

@@ 4 Experiments and Results
Dataset description: STARC-9 is a comprehensive multi-tissue classification dataset consisting of 630,000 high-resolution non-overlapping 256×256 pixel tiles extracted from 40x magnification (0.25 micrometers/pixel) WSI. It includes nine clinically relevant tissue types: ADI, LYM, MUS, FCT, MUC, NCS, BLD, TUM, and NOR, capturing diverse, fine-grained tissue morphologies. To facilitate rigorous validation of models trained on STARC-9, two independent validation sets were prepared. (i) STANFORD-CRC-HE-VAL-SMALL contains 18,000 tiles (2,000 per tissue type) obtained from 20 WSI separate from the cases used in STARC-9 and was used for preliminary model testing, yielding 79.59% and 85.9% accuracy for models trained on NCT-CRC-HE-100K (NCT) [Kather2019] and HMU-GC-HE-30K (HMU) [Lou2025GastricCancer], respectively. This highlighted the need for larger, more diverse training sets, as these models struggled with mixed tissue-type tiles, achieving only 60% overall per-WSI tissue mapping accuracy. (ii) The primary validation set, STANFORD-CRC-HE-VAL-LARGE, includes 54,000 tiles (6,000 per class) from 50 WSI independent from STARC-9 and STANFORD-CRC-HE-VAL-SMALL for performance evaluation. Training (STARC‑9) and validation (STANFORD-CRC-HE-VAL-SMALL and STANFORD-CRC-HE-VAL-LARGE) datasets were drawn from different patients, with no overlap between any of the datasets. Additionally, an external CURATED-TCGA-CRC-HE-VAL-20K dataset was prepared with 20,000 tiles extracted from 30 TCGA-CRC WSI to assess the robustness and generalizability of models trained on STARC-9. During benchmarking, models trained on STARC-9 (630,000 tiles, 9 classes), NCT (100,000 tiles, 9 classes), and HMU (30,000 tiles, 8 classes) were validated on seven overlapping tissue types (ADI, LYM, MUS, MUC, NCS, TUM, and NOR), ensuring fairness in performance evaluation.
@@
Model description: To evaluate the utility of STARC-9, we conducted a series of benchmarking experiments using a diverse set of deep learning models, including baseline CNNs, SOTA transformer models, and pathology-specific foundation models. The objective was to assess classification performance, generalizability, and practical utility compared to models trained (fine-tuned) on publicly available datasets like NCT and HMU. Baseline models included ResNet-50, EfficientNet-B7, KimiaNet, and ViT-base, while SOTA models included DeiT-B, Swin Transformer-Base, and ConvNeXT-Base. Pathology-specific foundation models such as CTransPath, HiPT, Prov-Gigapath, Path-DINO, CONCH, UNI, Virchow, and VIM4Path were also tested to assess their generalizability on diverse tissue morphologies. Each model was trained on the STARC-9, NCT, and HMU datasets with Macenko normalization [Macenko2009]. All models were fine-tuned with a batch size of 32, learning rate of 0.0001, weight decay of 1e-5, Adam optimization, and data augmentation (horizontal/vertical flips, random rotation, and color jittering) for 10 epochs. Maintaining identical training configurations, including batch size and optimizer settings, across datasets was important for a fair and unbiased model comparison. As our primary objective was to isolate the impact of the dataset on model performance, we kept the training configurations consistent to avoid possible confounding introduced by different hyperparameters being applied to the datasets. Model performance was evaluated using precision, recall, macro F1 score, accuracy, and the number of trainable parameters. STARC-9-trained models consistently outperformed models trained on other datasets, exhibiting better generalizability.
@@
Resource description: All experiments were conducted on the following platforms: (i) a local server with 8x NVIDIA L40S 48GB GPUs, and the Stanford (ii) Carina [Stanford_Carina] and (iii) Marlowe high-performance computing platforms [kapfer_2025_14751899].
@@
ResNet50 [He2016DeepResidual] 84.08 87.81 98.92 62.59 85.71 98.64 63.17 86.00 98.78 62.59 85.71 98.64 24 M
EfficientNet-B7 [Tan2019EfficientNet] 89.99 88.65 99.11 82.47 87.45 98.80 84.55 87.87 98.95 82.47 84.45 98.80 64 M
ViT-base [Dosovitskiy2021ViT] 92.71 91.57 98.49 84.25 90.29 98.09 87.30 90.87 98.28 84.25 90.29 98.09 86 M
@@
DeiT-B [pmlr-v139-touvron21a] 94.28 90.97 98.99 81.63 90.05 98.65 85.35 90.40 98.81 81.63 90.05 98.65 86 M
Swin Trans-base [Liu2022SwinTransformer] 90.11 93.17 99.09 79.05 91.88 98.80 82.52 92.46 98.94 79.05 91.88 98.79 87 M
KimiaNet [RIASATIAN2021102032] 87.25 88.60 99.03 71.53 86.67 98.72 71.53 87.04 98.87 68.69 86.67 98.72 7M
ConvNeXT-base [Liu2022ConvNet2020s] 91.95 92.09 99.01 82.82 91.07 98.36 85.56 91.50 98.68 82.82 91.07 98.36 88 M
@@
CTransPath [WANG2022102559] 90.11 93.17 99.34 79.05 91.88 99.00 82.52 92.46 99.16 79.05 91.88 99.00 87 M
HiPT [Chen2022ScalingViT] 90.92 93.21 98.64 74.51 91.99 98.32 77.41 92.54 98.47 74.51 91.99 98.32 86 M
ProvGigPath [Xu2024WholeSlideFoundationModel] 89.43 91.47 98.74 74.18 90.60 98.37 78.40 90.92 98.55 74.18 90.60 98.37 305 M
PathDino [Alfasly2024RotationAgnostic] 92.93 91.19 98.67 77.35 89.64 98.37 81.71 90.22 98.51 77.35 89.64 98.37 22 M
CONCH [Lu2024VisualLanguageCP] 91.53 91.41 98.56 75.69 90.02 98.19 78.08 90.52 98.37 75.69 90.02 98.19 87 M
UNI [Chen2024GeneralPurposeCP] 94.55 93.03 98.67 80.43 91.80 98.25 84.42 92.36 98.45 80.43 91.80 98.26 88 M
Virchow [Vorontsov2024FoundationModelCP] 92.51 92.35 98.63 79.02 91.23 98.28 82.05 91.69 98.45 79.02 91.23 98.28 305 M
@@ 4.1 Multi-Class Tissue Classification
Table 2 reports the precision, recall, F1‑macro, and accuracy metrics for the top‑performing models (with respect to accuracy in Table 1), when validated on STANFORD-CRC-HE-VAL-SMALL, STANFORD-CRC-HE-VAL-LARGE, and CURATED-TCGA-CRC-HE-VAL-20K. These top-performing models were: ViT‑Base trained on NCT, HiPT trained on HMU, and CTransPath trained on STARC‑9. For STANFORD-CRC-HE-VAL-SMALL, the STARC-9-trained CTransPath model achieved 99.75% precision, 99.73% recall, 99.74% F1- macro, and 99.73% accuracy - significantly higher than the other models, which showed lower recall and F1-macro scores. Similarly, on STARC-9-HE-VAL-LARGE, CTransPath maintained its lead, with 99.34% precision, 99.00% recall, 99.16% F1-macro, and 99.00% accuracy. Even on the more challenging STANFORD-TCGA-CRC-HE-20K set, CTransPath consistently achieved near-perfect precision (99.05%), recall (98.88%), and F1-macro (98.96%), demonstrating excellent generalization and robustness across diverse tissue types. To further evaluate the generalizability of the model trained on STARC-9, we curated a small test set of the seven common tissue classes taken from 10 WSI from the IMP‑CRS10K biopsy/polypectomy dataset [Neto2024IMPWSI]. In total, 1,093 image tiles were annotated for model performance validation, in which the STARC‑9-trained model achieved a 95.55% F1‑macro and 96.61% accuracy, consistently outperforming the HMU and NCT–trained models. It would also be interesting to evaluate the performance of the models trained on STARC-9 on the NCT and HMU datasets. However, as noted in Section 2 (and in reference [Ignatov2025]), the publicly available validation sets from NCT and HMU contain a substantial fraction of artifact‐laden or mislabeled tiles, as well as "ambiguous" tiles with more than one tissue type represented within the same tile, despite only a single tissue-type label being assigned to the tile. In order to utilize these two datasets as reliable validation datasets, pathologist re‑verification and correction/refinement of the tile‐level labels would be necessary, which is labor‐intensive and infeasible, given that the original WSI used to generate these two datasets were not publicly available for verification of label accuracy.
@@ 4.2 Tumor Tissue Segmentation
Among the most common downstream applications for multi-class tissue classification is tissue segmentation, especially tumor region segmentation, which allows for the automated identification and cropping of tumor-containing regions for subsequent annotation, ROI selection, diagnosis, and prognostication. This approach also facilitates downstream applications such as MSI [YAMASHITA2021132] and other biomarker status prediction, and supports survival modeling [Kather2019] for risk stratification and personalized treatment planning. In this context, we conducted experiments to evaluate the effectiveness of models trained on HMU, NCT, and STARC-9 for tumor segmentation, focusing on their ability to accurately identify tumor regions that are important for clinical decision-making. As there were no publicly available CRC WSI repositories with readily usable, high‑quality TUM masks for a larger scale experiment, and existing weakly‑supervised tools did not provide the precision required for generating tissue segmentation masks, we prepared a test set by enlisting pathologists to manually annotate (ground truth) the TUM region in patches of size 2048×2048 pixels using QuPath [Bankhead2017QuPath]. We selected 45 patches (3 per slide) from 15 Stanford WSI and 50 patches (2 per slide) from 25 TCGA-CRC WSI, which were fully independent of our training and validation sets, for a more controlled evaluation of model performance [Bokhorst2023DeepLearning]. Some patches contained mixed tissue types in order to evaluate the effectiveness of the trained models. For example, as shown in Figure 4, one region contained predominantly tumor, while the other included a mix of tumor and non-tumor tissue (NCS), providing more challenging segmentation.
@@ 5 Conclusion
Acknowledgments: Funding for this study was provided by the United States National Cancer Institute (NCI), National Institutes of Health (NIH) (R01 CA270437).
Technical Appendices
@@ Appendix D Additional experiments with recent models
We trained and evaluated recent models such as TransNeXt [shi2023transnext], OverLoCK [Lou11092403], and Beit-base [beit2022] with masked image modeling on the NCT, HMU, and STARC-9 datasets. The comprehensive results are presented in Table 4. While these recent models demonstrate competitive performance, they did not surpass our best-performing combinations: CTransPath on STARC-9, HiPT on HMU, and ViT-base on NCT (as reported in Table 1. Notably, models trained on STARC-9 consistently achieve superior performance across all architectures, validating our dataset’s quality and diversity.
@@
TransNeXt [shi2023transnext] 59.68 56.57 98.60 73.89 64.22 98.61 110M
OverLoCK [Lou11092403] 63.75 62.09 97.20 79.00 68.98 97.22 24.3M
Beit-base [beit2022] 59.95 76.48 98.17 74.08 86.17 98.20 86.5M
@@
TransNeXt [shi2023transnext] 59.62 55.05 98.58 73.55 62.78 95.68 110M
OverLoCK [Lou11092403] 63.69 55.56 98.85 79.29 64.09 98.87 24.3M
Beit-base [beit2022] 62.01 78.11 98.61 77.03 88.40 98.68 86.5M
@@
TransNeXt [shi2023transnext] 58.95 51.92 95.57 72.65 61.70 96.07 110M
OverLoCK [Lou11092403] 64.47 53.93 95.38 80.25 62.98 95.56 24.3M
Beit-base [beit2022] 57.45 72.57 97.93 71.75 82.11 97.87 86.5M
@@ Appendix G Ablation study
Number of samples per cluster ($m$): Determining the number of clusters ($K$) and $m$ is very challenging, and especially with histopathology-based tiles, it is difficult to set one value for the number of clusters and samples. We tested $m=100$ to $800$, finding that a small $(m\approx 100)$ yielded many tiny, redundant clusters, while a large $(m>800)$ resulted in mixing of tissue types. We found that $m\approx 400$ balanced intra-cluster purity and inter-cluster diversity. Table 6 shows the trade-off between the size of the cluster and the purity (inverse of the Shannon entropy) based on the fixed number of significant variations in TUM and NCS tissue morphology. As $m$ increases, purity decreases (entropy rises). We selected $m=400$, as it yielded substantial morphologic variation while maintaining low tissue-type admixture, representing the best balance for our goal. If deciding $m$ is complex for a dataset, it is recommended that the number of clusters and samples per cluster be set using $K=m=\sqrt{T}$ [sun2024pathgen16m16millionpathology], where $T$ is the number of tiles from a WSI.